Five years ago, Lilly's age was three times Cody's age. In two years, Lilly's age will be two times Cody's age. how old are Lilly and Cody now, in years?
Found 2 solutions by josgarithmetic, Boreal:
Answer by josgarithmetic(39617)
Answer by Boreal(15235)

Lilly is now x, Cody's is y
x-5=3(y-5)
x+2=2(y+2)
x=3y-10
x=2y+2
so 3y-10=2y+2
y=12 Cody
x=26 Lilly